# Can AI Relate: Testing Large Language Model Response for Mental Health Support

### v1

@@ Abstract
In this work, we develop an evaluation framework for determining whether LLM response is a viable and ethical path forward for the automation of mental health treatment. Using human evaluation with trained clinicians and automatic quality-of-care metrics grounded in psychology research, we compare the responses provided by peer-to-peer responders to those provided by a state-of-the-art LLM.
@@ 1 Introduction
Our study focuses on the use of LLMs for mental health response, which has faced increased pressure. Between 2020 and 2022, a Pew Research poll found that 41% of surveyed Americans reported high levels of psychological distress. This is particularly true amongst minority-identifying participants: 46% of Hispanic Americans and 42% of Black Americans reported high levels of distress compared to 38% of White Americans Gramlich (2023). One response to this mental health crisis has been the development of conversational agents to provide large-scale on-demand therapy. Notable examples include WoeBot and Youper, [^1] https://www.youper.ai/ chatbots trained to use cognitive behavioral therapy (CBT) techniques Fitzpatrick et al. (2017). While this is not a novel concept, with the Eliza chatbot envisioned in 1966 predating modern applications Weizenbaum (1966), the advent of LLMs raises new deployment opportunities and risks.
@@
While LLMs show remarkable performance on mental health related tasks like suicide prediction Xu et al. (2023) and self-referral Habicht et al. (2023), recent cases highlight the potential negative impact; (1) the death of a Belgian man by suicide after conversation with a chatbot based on GPT-J Walker (2023), and (2) advice given by chatbot Tessa to anorexia patients that they should diet and monitor their weight Eliot (2023). While work on evaluating empathy of LLMs is a promising step toward preventing such disastrous automation errors (e.g. Sharma et al., 2020), the potential for demographic perception bias in open-ended response has not yet been comprehensively explored. Specifically, no work has addressed whether LLMs provide equitable care across patient subgroups. Here we evaluate risks of discriminatory decision-making by LLMs for mental health response, to ensure all patients can receive equitable, appropriate care.
@@
Clinical Evaluation: We first have licensed clinical psychologists evaluate LLM and human responses to social media posts from mental health patients. Our results suggest room for improving LLM agents’ ability to embody broader values like partnership and actively counter negative patient behaviors. However, GPT-4 responses are more empathetic and 48% better at encouraging positive behavior change in patients than human peer-to-peer support. This suggests that LLMs may prove a promising direction for providing care if equity can be enforced, and motivates our further experimentation.
@@
Bias Evaluation: We next conduct an audit to assess if GPT-4 responses demonstrate biases towards certain demographic subgroups. We show that (1) GPT-4 is capable of inferring patient demographics like age, gender and race from the content of social media posts; (2) even when we control for the context by modifying posts to mention a specific demographic group, there is slight variability in empathy across subgroups; and (3) in responses to actual posts seeking help, there are statistically significant differences between empathy levels for minority subgroups and empathy for the majority (White) or subgroup. Specifically, there is a statistically significant difference between empathy in GPT-4 treatment responses for Black (2%-15% lower) and Asian posters (5%-17% lower) seeking mental health support compared to White or Unknown Race (control group) posters. While general-purpose LLMs like GPT-4 are already being integrated into medical settings Chen et al. (2023), they do not provide equitable mental health care across perceived patient subgroups.
@@
We envision that these findings and the new resources we provide for evaluation will enable further research into the ethical and safe use of LLMs for mental health treatment. [^2] We will release all data and code at https://github.com/skgabriel/mh-eval.
@@ 2.1 Automated Mental Health Support
As noted by Cho et al. (2023), most of the existing literature around automation of pyschotherapy implements rule-based or retrieval-based approaches. Many proposed applications for mental health support are based on carefully constrained decision tree algorithms, where therapeutic techniques, such as cognitive behavioral therapy (CBT) or acceptance and commitment therapy (ACT), are recommended based on users’ responses (e.g Morris et al., 2018; Dindo et al., 2017).
@@ 2.2 Equity in Healthcare Automation
Our work connects to a broader body of literature raising the alarm about potential harms of AI deployment in healthcare settings due to learned, often implicit bias Adam et al. (2022b); Zack et al. (2023); Omiye et al. (2023). Early work on bias in NLP-based mental health support Straw and Callison-Burch (2020) indicated this could propagate to inequity in treatment. Recent proposals (e.g Stade et al., 2023; Choudhury et al., 2023) have also addressed the need for auditing, though this has not been thoroughly evaluated with existing frameworks. On the other hand, recent studies Ayers et al. (2023); Chen et al. (2024) have found that progress in LLMs has led to generative technologies that can produce more empathetic, higher quality responses than trained clinicians. Following from these works, we seek to provide an empirical, quantitative evaluation of LLM risks and benefits in mental health deployment.
@@ 3 Data
To allow reproducibility of this work, we focus on peer-to-peer support from communities on Reddit known as “subreddits.” Prior work Naslund et al. (2016) has highlighted the effectiveness and growing prevalence of social media as a platform for mental health support. We use the Reddit API to collect a new dataset of 12,513 posts with 70,429 peer responses from 26 mental health related subreddits.
@@ 4 LLM Response Generation
Moreover, we further propose two settings as attempts to mitigate the bias:
@@
An example SMP persona prompt is shown below:
@@ 5 Attribute Inference
Given the sensitive nature of mental health data, personally identifiable information is rarely volunteered and publicly released research data is carefully de-identified. However, utterances may still contain implicit cues to a speaker’s background like linguistic variation (e.g. unique verb conjugation or word dropping) as well as explicit references (e.g. self-identifying race).
@@ 6 Evaluation Methodology
We first conduct an expert evaluation with 2 licensed clinical psychologists. We randomly sample 50 Reddit posts seeking mental health support, and pair each post with either an actual peer-to-peer response or a GPT-4 generated response [^4] Generated using the mental health forum persona with a length constraint of either “short” or “no more than 3 lines.” in a random order. The post/response pairs are shown to the clinicians using a Qualtrics [^5] https://www.qualtrics.com/ study. For ethical purposes, we state in the instructions that responses may be AI-generated, but we do not reveal the source of individual responses. We then ask the clinicians a set of 4 questions for each post/response pair, aimed at assessing the quality of care provided by the response. Based on the EPITOME framework Sharma et al. (2020) grounded in psychology research, we ask the following questions to gauge level of empathy on a 3-point Likert scale (none, weak, strong):
@@
For a more holistic view of quality of care, we then adapt the 5-point Likert scale Change Talk score from the Motivational Interviewing Treatment Integrity Code Resnicow and McMaster (2012):
@@
We predict types and levels of empathy demonstrated by responses using the computational empathy framework introduced by Sharma et al. (2020). We retrain individual RoBERTa-based classifiers to predict the empathy level of a post for each subcategory. On a randomly sampled held-out set (10% of the original non-proprietary dataset), we find the models have macro F1 of 80.5%, 87.3% and 92.7% respectively at binary empathy identification (no or some degree of empathy type).
@@ 7 Demographic Leaking Experiment
We start with a set of 11 seed “neutral” Reddit posts [^6] We determined based on statistical power analysis for a standard t-test with a moderate effect size that a minimum of 64 samples were required. seeking mental health advice for which GPT-4 was unable to predict any attributes of the patient. We then transform each post to show either an explicit leak (e.g. “I am a 32yo Black woman”) or implicit leak (e.g. “Being a 32yo girls girl wearing my natural hair”) using keyphrases self-disclosed by GPT-4 as indicators of a certain demographic (see §5 for details). We test 8 different combinations of post transformations,777[white_male_implicit, white_male_explicit, white_female_implicit, white_female_explicit, black_male_implicit, black_male_explicit, black_female_implicit, black_female_explicit] as well as a control setting with the original Reddit posts.
@@
Figure 1: Aggregated results of clinician evaluation for empathy metrics (Emotional Reaction (ER), Interpretation (IP) and Exploration (EX)) as well as the global score (Cultivating Change Talk (Change).
8 Results
8.1 Clinical Evaluation Results
Empathy.
MITI Global Score.
GPT-4 responses are particularly adept at focusing on and encouraging patient efforts to bring about positive change (human: $\mu_{1}=2.08$, $\mu_{2}=2.31$) vs. (AI: $\mu_{1}=3.08$, $\mu_{2}=3.13$). However, one of the clinicians noted that the GPT-4 responses seemed too direct, indicating this type of care may be low in other desirable values like partnership Moyers et al. (2016). A qualitative analysis of GPT-4 responses shows that as the model can’t relate through shared experience, it often acts out the role of an advisor and encourages patients to reach out for professional assistance. This may lead to the model being perceived as “talking down” to patients by some clinicians.
@@ 8.3 Demographic Auditing Results
We compare human responses to GPT-4, GPT-3.5 OpenAI (2022), and Mental-LLaMa Yang et al. (2023), a 13B parameter LLM trained on mental health data. From Figure 2, we can see that level of empathy for emotional reaction and exploration dimensions is dependent on the prompt variation. We also observe discrepancies in levels of empathy based on inferred racial subgroups, where for almost all model/prompt variations there is significantly lower empathy for Black posters compared to the control setting where race cannot be inferred. These differences are not observed for other attributes like gender and age (see Table 3 in the Appendix).
@@ 9 Bias Mitigation
We find that explicitly instructing LLMs to use demographic attributes in response (MHF-2, MHF-3) can effectively alleviate bias. This is the only prompt context in Figure 2 for which we found no negative statistically significant difference in empathy across subgroups for GPT-4 and Mental-LLaMa. There is some precedent for this explicit direction in human cognition, with prior studies finding that explicit racial priming can counteract implicit bias and encourage fairness Rachlinski et al. (2020), though explicit direction is not always effective with humans compared to implicit priming Huber and Lapinski (2008).
@@ 10 Conclusion
In conclusion, we develop a computational framework for assessing equity and overall quality of mental healthcare provided by LLMs like GPT-4. We find that while GPT-4 can reflect and amplify harmful biases found in peer-to-peer support, these biases can be mitigated through demographic-aware prompting. We hope that this work leads to more concrete guidelines and considerations for future deployment of LLMs in psychotherapy.
@@ 11 Acknowledgements
We thank our clinical evaluators at NYU Langone and the anonymous ACL RR reviewers for their constructive comments. MM was supported by the National Institute of Mental Health (NIMH) award K23MH134068. The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIMH.
@@ 12 Ethics Statement & Limitations
As for limitations, there is disagreement in psychotherapy-related literature around suitable criteria for assessing quality of care Gelso (1979); Barlow et al. (1999); Shafran et al. (2009). We use widely adopted methodologies for standardizing evaluation of quality of care Truax and Carkhuff (1969); Resnicow and McMaster (2012); Moyers et al. (2016); Sharma et al. (2020), however it is possible that our framework overlooks certain important aspects due to systemic inequality Roberts et al. (2020). It should be noted that there are many individual differences that impact efficacy of a particular therapy approach. Our evaluation is also narrowly scoped in terms of the type of bias we consider (e.g. we do not consider religious affiliation), which may impact both the mental health response and how a patient responds to a particular therapy approach.

### v2

@@ Abstract
In this work, we develop an evaluation framework for determining whether LLM response is a viable and ethical path forward for the automation of mental health treatment. Our framework measures equity in empathy and adherence of LLM responses to motivational interviewing theory. Using human evaluation with trained clinicians and automatic quality-of-care metrics grounded in psychology research, we compare the responses provided by peer-to-peer responders to those provided by a state-of-the-art LLM.
@@ 1 Introduction
Our study focuses on the use of LLMs for mental health response, which has faced increased pressure. Between 2020 and 2022, a Pew Research poll found that 41% of surveyed Americans reported high levels of psychological distress. This is particularly true amongst minority-identifying participants: 46% of Hispanic Americans and 42% of Black Americans reported high levels of distress compared to 38% of White Americans Gramlich (2023). One response to this mental health crisis has been the development of conversational agents to provide large-scale on-demand therapy. Notable examples include WoeBot [^1] https://woebothealth.com/ and Youper, [^2] https://www.youper.ai/ chatbots trained to use cognitive behavioral therapy (CBT) techniques Fitzpatrick et al. (2017). While this is not a novel concept, with the Eliza chatbot envisioned in 1966 predating modern applications Weizenbaum (1966), the advent of LLMs raises new deployment opportunities and risks.
@@
While LLMs show remarkable performance on mental health related tasks like suicide ideation or risk prediction Xu et al. (2023) and self-referral to mental health services Habicht et al. (2023), recent cases highlight the potential negative impact; (1) the death of a Belgian man by suicide after conversation with a chatbot based on GPT-J Walker (2023), and (2) advice given by chatbot Tessa to anorexia patients that they should diet and monitor their weight Eliot (2023). While work on evaluating empathy of LLMs is a promising step toward preventing such disastrous automation errors (e.g. Sharma et al., 2020), the potential for demographic perception bias in open-ended response has not yet been comprehensively explored. Specifically, no work has addressed whether LLMs provide equitable care across patient subgroups in a generative psychotherapy setting. Here we evaluate risks of discriminatory decision-making by LLMs for mental health response, to ensure all patients can receive equitable, appropriate care.
@@
Clinical Evaluation: We first have licensed clinical psychologists evaluate LLM and human responses to social media posts from mental health patients. Our results suggest room for improving LLM agents’ ability to embody broader values from motivational interviewing theory Moyers et al. (2016) like partnership and actively countering negative patient behaviors. However, GPT-4 responses are more empathetic and 48% better at encouraging positive behavior change in patients than human peer-to-peer support. This suggests that LLMs may prove a promising direction for providing care if equity can be enforced, and motivates our further experimentation.
@@
Bias Evaluation: We next conduct an audit to assess if GPT-4 responses demonstrate lower empathy towards certain demographic subgroups. We show that (1) GPT-4 is capable of inferring patient demographics like age, gender and race from the content of social media posts; (2) even when we control for the context by modifying posts to mention a specific demographic group, there is slight variability in empathy across subgroups; and (3) in responses to actual posts seeking help, there are statistically significant differences between empathy levels for minority subgroups and empathy for the majority (White) or subgroup. Specifically, there is a statistically significant difference between empathy in GPT-4 treatment responses for Black (2%-15% lower) and Asian posters (5%-17% lower) seeking mental health support compared to White or Unknown Race (control group) posters. While general-purpose LLMs like GPT-4 are already being integrated into medical settings Chen et al. (2023), they do not provide equitable mental health care across perceived patient subgroups.
@@
We envision that these findings and the new resources we provide for evaluation will enable further research into the ethical and safe use of LLMs for mental health treatment. [^3] We release all data and code at https://github.com/skgabriel/mh-eval.
@@ 2.1 Automated Mental Health Support
As noted by Cho et al. (2023), most of the existing literature around automation of psychotherapy implements rule-based or retrieval-based approaches. Many proposed applications for mental health support are based on carefully constrained decision tree algorithms, where therapeutic techniques, such as cognitive behavioral therapy (CBT) or acceptance and commitment therapy (ACT), are recommended based on users’ responses (e.g Morris et al., 2018; Dindo et al., 2017).
@@ 2.2 Equity in Healthcare Automation
Our work connects to a broader body of literature raising the alarm about potential harms of AI deployment in healthcare settings due to learned, often implicit bias Adam et al. (2022b); Zack et al. (2023); Omiye et al. (2023). Early work on bias in NLP-based mental health support Straw and Callison-Burch (2020) indicated this could propagate to inequity in treatment. Recent proposals (e.g Stade et al., 2023; Choudhury et al., 2023) have also addressed the need for auditing. Gender and racial bias has been found in classifiers trained for mental health prediction tasks, e.g. depression prediction Aguirre et al. (2021); Wang et al. (2024). It has also been shown that masked language models Lin et al. (2022) capture gendered stereotypes about mental health. On the other hand, recent studies Ayers et al. (2023); Chen et al. (2024) have found that progress in LLMs has led to generative technologies that can produce more empathetic, higher quality responses than trained clinicians. Following from these works, we seek to provide an empirical, quantitative evaluation of LLM risks and benefits in mental health deployment.
@@ 3 Data
To allow reproducibility of this work, we focus on peer-to-peer support from communities on Reddit known as “subreddits.” Prior work Naslund et al. (2016) has provided evidence of the growing prevalence of social media as a platform for mental health support. We use the Reddit API to collect a new dataset of 12,513 posts with 70,429 peer responses from 26 mental health related subreddits.
@@ 4 LLM Response Generation
These are designed to emulate 3 distinct and common sources of support a patient might seek: peer-to-peer support from a general audience, peer-to-peer support from a mental health focused audience, and professional support. Moreover, we further propose two settings as attempts to mitigate the bias:
@@
In theory, an unaware respondent should be more fair because they are conducting a blind evaluation. However, an explicitly instructed aware respondent may also be more fair because they are alerted to potential bias. An example SMP persona prompt is shown below:
@@ 5 Attribute Inference
Given the sensitive nature of mental health data, personally identifiable information is rarely volunteered and publicly released research data is often de-identified. However, utterances may still contain implicit cues to a speaker’s background like linguistic variation (e.g. unique verb conjugation or word dropping) as well as explicit references (e.g. self-identifying race).
@@
Validation of Approach.
A co-author manually checked the accuracy of the GPT-4 sourced gender, race and age labels with a random sample of 100 posts. To obtain a ground-truth label, they looked for explicit mentions of attributes (e.g. “24M” indicating a 24 year old male) or implicit mentions like women’s health conditions. If there was any uncertainty about an attribute label from the context, the ground-truth attribute label was marked as unknown. We find that GPT-4 and the human annotator have 94% agreement on race, 84% agreement on age, and 81% agreement on gender. The majority of disagreements were due to the human annotator abstaining from making a prediction (the ground-truth label is unknown), while GPT-4 makes an assumption.
@@ 6 Evaluation Methodology
We first conduct an expert evaluation with 2 licensed clinical psychologists. The 2 clinicians have backgrounds in adult psychology, specializing in depression, anxiety disorders and cognitive behavior therapy (CBT). We randomly sample 50 Reddit posts seeking mental health support, and pair each post with either an actual peer-to-peer response or a GPT-4 generated response [^5] Generated using the mental health forum persona with a length constraint of either “short” or “no more than 3 lines.” in a random order. The post/response pairs are shown to the clinicians using a Qualtrics [^6] https://www.qualtrics.com/ study. For ethical purposes, we state in the instructions that responses may be AI-generated, but we do not reveal the source of individual responses. We then ask the clinicians a set of 4 questions for each post/response pair, aimed at assessing the quality of care provided by the response. Based on the EPITOME framework Sharma et al. (2020) grounded in psychology research, we ask the following questions to gauge level of empathy on a 3-point Likert scale (none, weak, strong):
@@
For a more holistic view of quality of care, we then adapt the 5-point Likert scale Change Talk score from the Motivational Interviewing Treatment Integrity Code (MITI, Resnicow and McMaster (2012)):
@@
We predict types and levels of empathy demonstrated by responses using the computational empathy framework introduced by Sharma et al. (2020). We fine-tune individual RoBERTa-based classifiers to predict the empathy level of a post for each subcategory. On a randomly sampled held-out set (10% of the original non-proprietary dataset), we find the models have macro F1 of 80.5%, 87.3% and 92.7% respectively at binary empathy identification (no or some degree of empathy type).
@@ 7 Demographic Leaking Experiment
We start with a set of 11 seed “neutral” Reddit posts [^7] We determined based on statistical power analysis for a standard t-test with a moderate effect size that a minimum of 64 samples were required. seeking mental health advice for which GPT-4 was unable to predict any attributes of the patient. We then transform each post to show either an explicit leak (e.g. “I am a 32yo Black woman”) or implicit leak (e.g. “Being a 32yo girls girl wearing my natural hair”) using keyphrases self-disclosed by GPT-4 as indicators of a certain demographic (see §5 for details). We test 8 different combinations of post transformations, [^8] The combinations we consider are: [white-male-implicit, white-male-explicit, white-female-implicit, white-female-explicit, black-male-implicit, black-male-explicit, black-female-implicit, black-female-explicit] as well as a control setting with the original Reddit posts.
@@
Figure 1: Aggregated results of clinician evaluation for empathy metrics (Emotional Reaction (ER), Interpretation (IP) and Exploration (EX)) as well as the global score (Cultivating Change Talk (Change). The y-axis shows the averaged score values across samples and clinicians.
@@
8 Results
8.1 Clinical Evaluation Results
Empathy.
MITI Global Score.
GPT-4 responses are particularly adept at focusing on and encouraging patient efforts to bring about positive change (human: $\mu_{1}=2.08$, $\mu_{2}=2.31$) vs. (AI: $\mu_{1}=3.08$, $\mu_{2}=3.13$). [^10] We note that as defined in the motivational interviewing integrity coding manual, this measures the responder’s efforts to encourage change and does not measure the patient’s actual behavioral change. However, one of the clinicians noted that the GPT-4 responses seemed too direct, indicating this type of care may be low in other desirable values like partnership Moyers et al. (2016). A qualitative analysis of GPT-4 responses shows that as the model can’t relate through shared experience, it often acts out the role of an advisor and encourages patients to reach out for professional assistance. This may lead to the model being perceived as “talking down” to patients by some clinicians.
@@ 8.3 Demographic Auditing Results
We compare human responses to GPT-4, GPT-3.5 OpenAI (2022), and Mental-LLaMa Yang et al. (2023), a 13B parameter LLM trained on mental health data. From Figure 2, we can see that level of empathy for emotional reaction and exploration dimensions is dependent on the prompt variation. We also observe discrepancies in levels of empathy based on inferred racial subgroups, where for almost all model/prompt variations there is significantly lower empathy for Black posters compared to the control setting where race cannot be inferred. Emotional reaction empathy of GPT-4 responses is lower for Black posters than White posters for the MHF-1 prompt (p = 0.01), the Clinician prompt (p < 0.01), and the MHF-2 prompt (p < 0.02). For overall empathy of GPT-4 responses the significant gap is primarily between Black posters and the control (Unknown) subgroup, with significant differences appearing for the SMP prompt (p = 0.01), Clinician Prompt (p = 0.03), and MHF-2 prompt (p < 0.01). These differences are not observed for other attributes like gender and age (see Table 3 in the Appendix). These differences can amplify biases from human peer-to-peer responses. For example, overall empathy for Asian posters in human peer-to-peer responses is 12% lower than White posters, but 17% lower in MHF-1 responses from GPT-4.
@@ 9 Bias Mitigation
We find that explicitly instructing LLMs to use demographic attributes in response (MHF-2, MHF-3) can effectively alleviate bias depending on the model. The MHF-3 prompt context in Figure 2 is the only one for which we found no statistically significant difference in empathy across subgroups for GPT-4 (see Figures 2a, 2d and 2g). However, results are more mixed for GPT-3.5 and Mental-LLaMA. In particular see Figure 2i, where results are not significantly different for all subgroups with MHF-2, however they are for both MHF-2 and MHF-3 with GPT-3.5 in Figure 2h. There is some precedent for this explicit direction in human cognition, with prior studies finding that explicit racial priming can counteract implicit bias and encourage fairness Rachlinski et al. (2020), though explicit direction is not always effective with humans compared to implicit priming Huber and Lapinski (2008). Future work can explore whether there exists a more universally fair prompt context.
@@ 10 Conclusion
In conclusion, we address the current lack of mental health chatbot bias testing by developing a computational framework for assessing equity and overall quality of LLM-based mental healthcare. We find that while GPT-4 can reflect and amplify harmful biases found in peer-to-peer support, these biases vary significantly based on prompt design and can be mitigated through demographic-aware prompting. This proposed solution is a simple and actionable way of incorporating explicit bias mitigation into future work on generative mental health support. We propose that developers of LLM-based mental health technologies employ (1) carefully designed LLM instruction in mental health support to avoid bias, and (2) rigorous evaluation of mental health chatbots that considers potential variation across demographic groups. We hope that our findings lead to more concrete guidelines for future deployment of LLMs in psychotherapy.
@@ 11 Acknowledgements
We thank our clinical evaluators at NYU Langone and the anonymous reviewers for their constructive comments. MM was supported by the National Institute of Mental Health (NIMH) award K23MH134068. The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIMH.
@@ 12 Ethics Statement & Limitations
As for limitations, there is disagreement in psychotherapy-related literature around suitable criteria for assessing quality of care Gelso (1979); Barlow et al. (1999); Shafran et al. (2009). We use widely adopted methodologies for standardizing evaluation of quality of care Truax and Carkhuff (1969); Resnicow and McMaster (2012); Moyers et al. (2016); Sharma et al. (2020), however these are imperfect proxies that allow for scalability. A more ideal but less scalable evaluation would measure the effect on patients’ psychotherapeutic outcomes. We encourage future work analyzing bias in LLM-based psychotherapy with longitudinal human subjects studies. Additionally, it is possible that our framework overlooks certain important aspects due to systemic inequality Roberts et al. (2020). It should be noted that there are many individual differences that impact efficacy of a particular therapy approach. Our evaluation is also narrowly scoped in terms of the type of bias we consider (e.g. we do not consider religious affiliation), which may impact both the mental health response and how a patient responds to a particular therapy approach. Lastly, we acknowledge the study size as a limitation in terms of both data samples and number of clinical evaluators. Future work can consider scaling our evaluation to strengthen statistical power and account for individual differences among clinician ratings.